Shift report 20010818 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Aug 18 2001 - 16:17:44 EDT

  • Next message: Apache: "Shift report 20010819 16:00-24:00"

    ------------------------------------------------------------
    Date: Sat Aug 18 16:17:44 2001, Shift: 08:00-16:00
    Supervisor: JIJ, with: TL
    Log book#107 pages 121 - 123
    
    Shift Summary:
    ------------------------------------------------------------
    1. A couple of quench link interlocks when we had rather good beam, at 10:36 and 13:33
       MCR called and reported "cfe-2c-brahms response failure". Related to RHIC VME crate for magnet controls, i.e. RHIC's business.
       MCR rebooted the crate and things look fine.
    
    2. Run 4268: 67.9k trigger 1 events
       Run 4271: 65.5k trigger 1 events
       (That sums up to ~800k trigger 1 events at B polarity setting started 8/16 i.e. ~200k to go.) 
    
    3. T3 anodes 2 are down. Probably broken wire
       One bad chip in FEE card on T1 row 4. Should be swapped at next access.
       	
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Aug 18 2001 - 16:18:36 EDT